Find (f. g)(x). g(x) = x + 7 (r. g)(x)
Answer #1

① given f(x) = 42, g(x)=x+7. (f. g)(x) = Assign the function f = 42 => (42.9()) () Anign the function 9=x+7. -) 44(x+7) = 4x²